OpenStudy (anonymous):

the circular blade on a saw has a diameter of 7.5 inches and rotates at 2400 revolutions per minute. A) find the angular speed in radians per sec B) Find the linear speed of the saw teeth ( in feet per sec) as the contact the wood being cut. Formulas--> Linear speed= arc length/time= s/t Angular speed= central angle/time arc length s= radius r * central angle(in radians)

OpenStudy (anonymous):

the answer in the book is 80pi rad per sec for a and 25pi ft per sec for b but i dont know how to get to that...

OpenStudy (anonymous):

There are 2pi for every revolution

OpenStudy (anonymous):

so we multipy 2pi * 2400 this is angular speed in minutes, we want is in sec ; so divide by 60 2pi * 2400/60 =80pi

OpenStudy (anonymous):

linear speed = radius time angular speed

OpenStudy (anonymous):

radius= 7.5/2=3.75 inches=.3125 feet

OpenStudy (anonymous):

b)linear speed = radius x angular speed .3125 x 80pi=25 pi

OpenStudy (anonymous):

ur wrong you wouldn't divide by 60 you would multiply thus getting the wrong answer

OpenStudy (anonymous):

for the min to sec thing

OpenStudy (anonymous):

but it matches your answer

OpenStudy (anonymous):

the angular speed is wrong you would multiply 60 to convert min to sec

OpenStudy (anonymous):

If I run 60 miles per minutes, how many miles do I run in a second?

OpenStudy (anonymous):

1 mi per sec?

OpenStudy (anonymous):

yeah, did you divide or multiply by 60?

OpenStudy (anonymous):

divide

OpenStudy (amistre64):

2400 1 min -----*------ = 240/6 rps 1 min 60 sec

OpenStudy (anonymous):

okay i got what your saying thanks :) and im sorry that i said you were wrong

OpenStudy (anonymous):

no problem, glad to help